# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v4

@@ 5 Results
Figure 7: Thermodynamic & pattern-matching efficiency. In (a)-(d) the thermodynamic efficiency, $\eta$, as a function of $r$. The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i)-(l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant "length" scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure. In (e)-(h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In panels (b)-(f) and (d)-(h) a noticeable cohesion of the peaks of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. The efficiency peaks, both $\eta$ and $\theta$, are always in the amorphous region of the phase plane. For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star-graph to the amorphous region. This agrees with the argument that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
Figure 8: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a)-(d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Figure 7. $\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i)-(l) of Figure 7. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
5.1 Performance at the emergence of scale-invariance
In this section, we discuss the relation between the location of the pattern matching efficiency, $\theta$, on the phase plane of the algorithm and the complexity of the dataset, see Figure 2. We turn our attention to the panels (e)-(h) and (i)-(l) in Figure 7 and panels (a)-(d) in Figure 8.
For the CIFAR10 and FashionMNIST datasets, there is a rough correspondence between the $r$-value where the structural transition emerges, $r\sim 0.5$ for CIFAR10 and $r\sim 0.35$ for FashionMNIST, and the $r$-range where the peaks of each of the two $\theta$ efficiencies are observed, $r\in(0.30,0.50)$ and $r\in(0.35,0.55)$ respectively. Note that for the FashionMNIST, also the thermodynamic $\eta$ peaks in the same range of $r$.
On the other hand, for the MNIST and NIST datasets, $\theta$ peaks in ranges of $r$ which are farther away from the respective bifurcation points $r\sim 0.4$ and $r\sim 0.5$. The term "far away" is quantified by the purple-shaded areas in panels (e)-(h) of Figure 7. Note the correspondence of the peaks of $\eta$ and $\theta$ for the two datasets in terms of $r$. The latter observation on the correspondence of $\theta$ and $\eta$ also holds for the MNIST, if we include $\theta$ points in the one standard deviation illustrated by the shaded-purple area.
To explain the varying locations of peak efficiencies, whether near the transition border or farther from it, we refer to an argument in the literature [37, 10, 38]. A system needs tuning, or self-tuning, at the edge of a phase transition only when the task demands a high level of complexity, specifically, this was investigated in the context of second-order phase transitions [37, 10].
For the more complex datasets CIFAR10 and FashionMNIST, the optimal $\theta$ values are closer to the transition border. In contrast, the optimal values for the simpler NIST and MNIST datasets are located farther from their respective transition points. This observation also holds for the thermodynamic $\eta$ for the FashionMNIST, MNIST, and NIST. Moreover, the MNIST, an intermediate dataset in terms of complexity, is an interesting case. It shows two maxima for the thermodynamic efficiency, $\eta$. The global maximum in Figure 7(c) is far away from the critical value of $r$ illustrated by the bifurcation in panel (c) of Figure 8, while the local maximum is at $r\sim 0.35-0.40$, which corresponds to the critical value of $r$ indicated by the bifurcation.
These observations align with the hypothesis that the system exploits its proximity to the phase transition boundary only when presented with a relatively complex task. The overall discussion highlights the importance of the dataset under analysis, particularly its disordered distribution of points within a high-dimensional feature space.
@@ 5.1 Performance at the emergence of scale-invariance
Finally, we observe that while the value of $r$ where a structural transition appears is dependent on the dataset, i.e. $r=0.50$ for the NIST, $r=0.40$ for the MNIST, $r=0.35$ for the FashionMNIST and $r=0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is roughly the same for all the four datasets. Interestingly, the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplementary Section S2). We speculate about the possibility that this is due to the constraint on sparsity as it is an invariant feature over the cases investigated in this work.
6 Discussion
@@
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [39]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an efficient internal representation of the complex world - or the dataset - they are posed in contact with. Such internal representation can be encoded in terms of abstract probability distributions [9, 10] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [40], or in a mixed way as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine to achieve statistical generalization [41].
@@
When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli possibly drawn from the same distribution they adapted to [42]. Moreover, the same system needs to cope with the cost of sending information across its structure which can cause an enhanced exploitation of resources. In the context of machine learning, the cost can be quantified by the number of computations to be done before the satisfying answer to the task can be reported, e.g. class-matching.
The recently developed theoretical tools of statistical physics for networks [5, 4, 43, 22, 3, 1] permit to quantify the diversity in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As fastness of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex networked structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two, and a variational principle in terms of thermodynamic efficiency can be devised to explain why a certain structure emerge in place of an other [1].
@@
In our work, this variational principle is computationally employed to explain the structural properties of the network subclass composed of sparse heterogeneous hierarchical trees produced by the Portegys protocol that, similar to living systems, need to learn the diversity of the environment they are in contact with and consequently self-organize their structure. The tree formation undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
@@
Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets with different levels of complexity. The optimal pattern-matching efficiency was always found in the amorphous regime, Nevertheless, when the dataset was more difficult, an optimal efficiency to the pattern-searching task was found close to the transition border separating the star-graph and the scale-invariant regimes.
@@
Moreover, empirical evidence suggests that the thermodynamic efficiency and the pattern-matching efficiency show some consistency, namely when we look at the global maxima of the two in the $r$-range. Even though the two processes involved possess significant differences, namely the Portegys search phase is a directed process starting from the root node, while the diffusion propagator describes an ergodic process propagating from any possible node, we think that is reasonable to expect some similar behavior of the two efficiencies, as both quantities, in the end, estimate a trade-off between the flow of a process and the encoding capabilities of the system in its network structural traits when confronted to the disorder embedded in a dataset of real-world patterns.
@@
In both cases, the "flow" is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to convince ourselves that these two quantities describe the flow in the network as the former counts the number of visited nodes across the traversing of the network by the Portegys search process, while the latter is minus the logarithm of $Z$, with $Z$ proportional to the average return probability after time $\tau$ of a continuous time edge-centric random walker [4]. They are not exactly alike as only the latter is an ergodic process; still, they can be used as estimates of the "flow".
@@
We have discussed the denominator of the two efficiencies: $\eta$ and $\theta$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se as evidenced by the fact that the accuracy is not necessarily a monotonic decreasing function of the control parameter $r$, but it can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.

### v5

@@ III.3 Phase portrait
Figure 7: Thermodynamic & pattern-matching efficiency. In (a–d) the thermodynamic efficiency, $\eta$, as a function of $r$. The thermodynamic efficiency is measured in the $\tau$-range of one order of magnitude from the bifurcation of the specific heat, $C$, depicted in panels (i–l). The $\tau$-resolution scale $\tau\sim 10$ identifies the relevant “length” scale resolving the transition from the single-peak to the emergence of a roughly plateau region, $C=C_{0}>0$, reflecting a roughly scale-invariant structure. In (e–h) the pattern-matching efficiency, $\theta$, is plotted as a function of $r$ for the four investigated datasets of increasing complexity (see arrow at the bottom of the figure). The shaded purple area indicates the tolerance $r$-range for the optimal pattern-matching efficiency, which includes points in one standard deviation from the global maximum indicated by the vertical red-dashed line. In panels (b),(f) and (d),(h) a noticeable cohesion of the peaks of the two efficiencies is observed. Vertical dashed lines indicate the global maxima of the efficiencies. The efficiency peaks, both $\eta$ and $\theta$, are always in the amorphous region of the phase plane. For datasets of higher complexity, they tend to be closer to the structural phase transition border separating the star graph from the amorphous region. This agrees with the argument that a fairly simple dataset does not benefit from the vicinity of the phase transition. Quantities are averaged over 50 independent simulations.
III.4 Efficiency
In Fig. 7, panels (a–d) depict the thermodynamic efficiency, $\eta$, and panels (e–h) depict the pattern-matching efficiency, $\theta$. See Equations 5 and 6 respectively.
The thermodynamic efficiency [1], $\eta=(Q+W)/W$, quantifies the trade-off between two quantities. The former, $Q=-\delta S/\tau$, is proportional to the variation of entropy content, $\delta S$, from the case of isolated nodes (see Methods). Thus, it estimates the ability to express diversity to a stochastic perturbation, or in other terms, the internal representation capabilities provided by the network topological traits. The latter is the variation of information flow with respect to a network of isolated nodes, $W=\delta F$; so to speak, how fast information travels across the network (see Methods).
The pattern-matching efficiency, $\theta$, defined in Equation 6, quantifies the trade-off between accuracy and number of visited nodes in the pattern-recognition task, representing the functional aspect of the algorithm (see Methods).
The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in panels (e–h) indicate a $r$-domain tolerance for $\theta$. The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
The two efficiencies $\theta$ and $\eta$ show some consistency in the x-range of highly efficient structures, see panels (b),(d) related to the NIST and the FashionMNIST. If we include points in the tolerance $r$-range spanned by the purple-shaded areas in panels (e–h), we observe the same consistency for all datasets when $\eta$ is computed at $\tau=10$ (red dashed lines in panels (a–d)). This consistency in the maxima suggests that the two efficiencies might be related, in the specific context of the algorithm here investigated. In the Sec. IV we discuss this matter.
The specific resolution scale, $\tau\sim 10$, at which we compute $\eta$ is the scale involved in the mesoscale structural transition from star graph to amorphous scale-invariant, that we discussed in Sec. III.3 and Fig. 6. Such critical resolution scale is identified by the bifurcation of the specific heat peaks in Fig. 7(i–l) and in Fig. 8(a–d).
In the seminal work in Ref. [1], it is suggested to use as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales where to compute the thermodynamic efficiency, $\eta$. The Fiedler eigenvalue provides the largest network time scale where the diffusion process has reached equilibrium. Nevertheless, $\tau_{diff}$ seems to only approximate the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$, especially in the star-graph region for small $r$. Panels (a–d) of Fig. 8 depict the $\tau_{diff}$ curves in red, next to the $\tau^{(1,2)}_{peaks}$ curves. The key difference is the thermal ensemble average: a weighted average, that employs all the eigenvalues to compute the specific heat and identify its peaks.
Differently from what was proposed in Ref. [1], we don’t use $\tau_{diff}$, or $\tau_{peak}^{(2)}$ in Fig. 8. Because that would mean adopting the largest possible resolution, which would not monitor the structural mesoscale changes involved in the transition, from the single trivial scale of the star graph to the more complex topological traits of the amorphous networks in the scale-invariant regime.
The resolution scale, $\tau\sim 10$, is shared across all topologies in the phase plane and, remarkably, defines an “ultra-violet” cut-off in the characteristic frequency of the networks across the transformation from one phase to the other. This is in analogy to the process of integrating out the large eigenvalues responsible for the finer details, as discussed in the context of the Laplacian-Renormalization group [35]. Note in Fig. S1 of the Supplemental Material [63] that our choice of $\tau=1/\lambda\sim 10$ is supported by the exclusion of the resolution region where the power-law behavior of the spectral density, $\mathcal{P}(\lambda)$, is lost. In Sec. S3 of the Supplemental Material [63], the efficiency, $\eta$, is computed in the range interval $\tau<10$. Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching efficiency.
@@ III.4 Efficiency
Moreover, we observe that, while the value of $r$ where a structural transition appears is dependent on the dataset, i.e. $r=0.40$ for the NIST, $r\simeq 0.40$ for the MNIST, $r\simeq 0.35$ for the FashionMNIST, and $r\simeq 0.50$ for the CIFAR10, the $\tau\sim 10^{1}$ scale is the same for all the investigated datasets. Finally, we note that the bifurcation $\tau$-scale is also unaffected by the number of nodes in the network (see Supplemental Material [63] Sec. S2).
Figure 8: Role of the ensemble average on the estimation of the $\tau$-scale. Panels (a–d) depict the $\tau^{(1,2)}_{peak}$ and $\tau_{diff}$ as a function of $r$. $\tau^{(1,2)}_{peak}$ are the $\tau$-location of the first and last peak of the specific heat in Fig. 7. $\tau_{diff}$ is the inverse of the Fiedler eigenvalue which provides a good estimation of the end of the diffusion process only for sufficiently large $r$. Note that the red curves retrace the profile of the upper branch of the specific heat for $\tau\gtrsim 10$ in panels (i–l) of Fig. 7. At smaller $r$ values the equilibrium time is better reflected by $\tau^{(2)}_{peak}$. The equilibrium of the diffusion process is an increasing function of $r$. $\tau_{diff}$ curves are averaged over 50 independent simulations. The $\tau^{(1,2)}_{peak}$ curves are computed from the specific heat curves which are averaged over 50 independent simulations.
III.5 Datasets’ phase plane
The topological complexity of the networks increases at a critical $r$, from exhibiting the trivial traits of a star graph to the structure of the amorphous network with a roughly power-law spectrum. Thus, in the phase plane, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau between the first and the last peak. This is illustrated in Fig. 7(i–l) for each investigated dataset. Such transition is also depicted in Fig. 8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in Fig. 7(i–l). See also Sec. S4 in the Supplemental Material [63] which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
The MNIST, an intermediate dataset in terms of complexity, is an interesting case. We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see Fig. 7(k)). This is an intrinsic characteristic of the dataset, as this feature is only mitigated for larger networks, as shown in Figs. S2 and S7 of the Supplemental Material [63].
This intermediate specific heat peak is reflected on two maxima for the thermodynamic efficiency, $\eta$, and a two-step transition in the amorphous region, in Fig. 7(c),(k). The local maximum is at $r\in[0.35-0.45]$, depending on the specific $\tau$-scale. Such a range aligns with the transition at $r\simeq 0.40$ indicated by the bifurcation of the two specific heat peaks in Fig. 8. The global maximum in Fig. 7(c) is close to the second-step transition at $r\simeq 0.6$. Although the local maximum in $\eta$ is not detected by $\theta$ in Fig. 7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area, used to compare $\eta$ and $\theta$ peaks in terms of $r$. This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Fig. 7(c).
Finally, note in Fig. 8 the slow down of diffusion due to the increasing “volume” of the network discussed in Sec. III.3 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$. We remark that $\tau^{(2)}_{peak}$ is a function of solely the parameter $r$. The latter shapes solely the degree distribution, as the number of nodes and edges is fixed across the phase plane and fixed in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.
III.6 Efficient performance at the emergence of scale-invariance
In this and the next section, we discuss the relation between the location of the maximum pattern-matching efficiency on the algorithm’s phase plane (Fig. 7(e–h)) and the dataset’s complexity (Fig. 3).
All optimal networks are found in the amorphous region of the phase plane where the dilation symmetry is observed in the spectral density and reflected in the specific heat plateau. To discuss this observation we refer to the following argument [67, 11, 68]: a system needs tuning, or self-tuning, at the edge of a phase transition when the task demands a high level of complexity. Specifically, this was investigated in the context of second-order phase transitions [67, 11]. But what is the advantage of being close to a second-order phase transition? One possible advantage comes from the power-law behavior of the density function of some quantities, e.g. density of cluster sizes in percolation, or density of avalanche sizes and duration in systems with temporal dynamics. The power law provides the system with a large range of availability from the density of the quantity of interest, and the system may exploit such availability as a resource in the task, whether it is adaptation to a complex environment or other goal-oriented ones.
In our case, the amorphous region of the phase plane, where the network traits display self-similar structures in a limited $\tau$-resolution range, exhibits the power law of the spectral density, governing the network frequencies. Thus, better performances in terms of efficiency, particularly pattern-matching efficiency, can be explained in terms of network frequencies, which provide effective resources for the real-world disorder structuring task.
III.7 Spectral dimension estimates the dataset complexity
We now examine the location of the optimal $\theta$ efficiency in the amorphous region of the phase plane for each dataset. Thus, for each dataset, we inspect the traits of the highly efficient structures. We resort to the spectral dimension [39, 40, 41], $d_{s}$, which generalizes the concept of Euclidean dimension to disordered structures. It takes non-integer values between 1 and 3 and the same value of the Euclidean dimension in the case of d-dimensional lattices. The spectral dimension gauges the effective large-scale connectedness of disordered structures, with high values of $d_{s}$ indicating high topological connectedness, and is related to the asymptotic behavior of diffusion [40], but also to electrical transport and thermal stability of folded proteins [42]
Depending on where the maximal $\theta$-efficiency is observed in the amorphous region of the phase plane in Fig. 7, we see networks changing the value of the specific heat plateau, reflecting the spectral dimension, $d_{s}=2C_{0}$ [34], as we discussed in Sec. II. We also observe that the specific heat plateau is an approximately decreasing function of $r$ in the scale-invariant region of the phase plane in Fig. 5(a). See also Sec. S4 in Supplemental Material [63].
For complex datasets, such as the CIFAR10 and FashionMNIST, the optimal $\theta$s are closer to the transition border delimiting the star graph from the amorphous phase. In contrast, for the simpler NIST and MNIST datasets, the optimal values are farther from their respective transition points. This observation also holds for the thermodynamic efficiency, $\eta$, noted the correspondence of the peaks of $\eta(\tau=10)$ and $\theta$ for these datasets in terms of $r$. A special case is again the MNIST, with its two-step transition in the amorphous region, where local maximum $\eta$ are at the first transition, $r\simeq 0.35$, and the global maxima of $\theta$ and $\eta$ are at $r\simeq 0.6$.
We measure the spectral dimension for each dataset (N=2000) by averaging over the exponents of the spectral density $\mathcal{P}(\lambda)\sim\lambda^{\gamma}$. The exponents are obtained fitting the power law in the range $\lambda\in[1\times 10^{-3},2\times 10^{-1}]$ for all those networks which better adapted to the dataset, i.e. those included in the purple $r$-range of maximal $\theta$-efficiency in Fig. 7(e-h). The spectral dimension, $d_{s}=2(\gamma+1)$ [41], is plotted with red dots for each dataset in Fig. 9. We removed the value $\gamma(r=0.35)$ when measuring the spectral dimension of the CIFAR10 because the spectral distribution was too close to the transition to obtain a good fit of the power law.
We also estimated the spectral dimension by measuring the specific heat plateau [35, 36], $C_{0}$, over the resolution range $\tau\in[1,2]\times 10^{1}$. The measurements, $d_{s}=2C_{0}$, are shown with blue triangles in Fig. 9.
Fig. 9 shows that the amorphous networks’ spectral dimension is a good estimator of the dataset complexity in Fig. 3. For higher-complexity datasets (ranked from left to right, from the low complexity NIST to the high complexity CIFAR10) we have increasing spectral dimension, $d_{s}$. In other words, the spectral dimension provides a way to gauge the complexity of the dataset after applying the tree generation protocol. Moreover, we can assess that high-complexity datasets are arranged in topological structures that are more compact when compared to those of lower complexity.
Figure 9: Datasets’ spectral dimension. The amorphous networks’ spectral dimension is a good estimator of the dataset complexity (Fig. 3). The red points indicate the spectral dimension [41], $d_{s}$, measured from the power-law exponent of the spectral density. The blue dots are obtained by monitoring the specific heat plateau [35, 36], $d_{s}=2C_{0}$, averaged over the resolution $\tau\in[1,2]\times 10^{1}$. All points are averaged over the $r$-range of the phase plane of maximal $\theta$-efficiency, see the shaded purple areas in Fig. 7(e–h), and over 50 independent simulations.
IV Discussion
@@
The ability of a system to survive is the capability of concentrating a flow of order from the environment - “drinking order from the environment” [69]. In other words, systems, whether living systems or machine(-learning) systems, need to learn an internal representation of the complex world - or the dataset - they are posed in contact with. Ideally, to be competitive, such internal representation should be efficient in terms of desired function and available resources. It can be encoded in terms of abstract probability distributions [10, 11] or into the very network features of the system, as it happens in genetic regulatory networks playing a central role in morphogenesis [13]. Or in a mixed way, as in artificial neural networks where the engineer designs a complicated distribution through the definition of the architecture and the training phase fills in the parameters to obtain the desired function-approximation machine able to achieve statistical generalization [70].
@@
When confronted with the real-world environment, to survive, living systems not only need the capability of interpreting the environment, but they also need the capability to quickly react with a coherent response to stimuli, possibly drawn from the same distribution they adapted to [71]. Moreover, the same system needs to cope with the cost of sending information across its structure, which can lead to an enhanced exploitation of resources. In machine learning, the cost can be quantified by the number of computations required before the satisfying answer to a task, such as class matching, can be reported.
@@
Here, we used an algorithm [27] that creates and navigates trees of real-world patterns to investigate the intertwined role of structure and function in hierarchical self-organized tree systems in a simple setting. The algorithm, akin to living systems, adapts to the diversity of the environment it is in contact with, represented here by the real-world dataset. It self-organizes structuring the disordered collection of high dimensional vectors into its network traits. The tree formation process undergoes two phase transitions at two different scales: the former is a mesoscale transition with the emergence of an approximately scale-invariant structural behavior progressively involving larger scales; the latter is a large-scale transition from the irregular scale-invariant structure to the regular-scale invariant 1d lattice.
@@
Given that the original purpose of the algorithm was an efficient pattern-searching method, we have investigated the pattern-matching efficiency and its positioning on the phase-plane for four different datasets exhibiting distinct complexity levels.
The recently developed theoretical tools of statistical physics for networks [33, 54, 72, 34, 35, 1] permit to quantify the diversity expressed in the traits of the network, with the von Neumann entropy, and the transmission flow of information, with the Helmholtz free energy. As the speed of information flow and sufficiently rich representation of the external world are two competing capabilities, the formation process of a complex structure can be interpreted as the macroscopic thermodynamic outcome of the delicate balance between the two. Thus, a variational principle gauged by the thermodynamic efficiency can be devised to explain why a certain structure emerges in place of another [1]. In our work, this variational principle and the other tools provided by the theory are computationally employed to monitor the structural properties of sparse heterogeneous hierarchical trees produced by the Portegys protocol.
@@
Empirical evidence suggests that the thermodynamic and the pattern-matching efficiencies show some consistency, namely when we look at the global maxima of the two in the $r$-range. These two quantities involve two distinct processes that possess significant differences. On the one hand, the Portegys search phase is a directed process starting from the root node that navigates the structure under a homophily principle. On the other hand, the diffusion propagator describes an undirected ergodic process propagating from any possible node. Moreover, while the thermodynamic efficiency is defined based on general statistical physics arguments, and thus it is task-agnostic; the pattern-matching efficiency is defined based on the specific task under consideration.
@@
Still, we think it is reasonable to expect a similar behavior of the two efficiencies. Both quantities estimate a trade-off between the flow of a process and the encoding capabilities of the system’s topological traits. In both cases, the “flow” is quantified by a monotonic decreasing quantity as a function of $r$. For the Portegys we have the logarithmic number of computed distances, while for the Laplacian process, we have $\delta F=F+\frac{\log N}{\tau}$, where $F=-\frac{\log Z}{\tau}$. It is easy to see that these two quantities describe the flow in the network. The former counts the number of visited nodes across the traversing of the network by the Portegys search process. The latter is minus the logarithm of $Z$, with $Z\sim\mathcal{R}(\tau)$ where $\mathcal{R}(\tau)$ is the average return probability after time $\tau$ of a continuous time edge-centric random walker [54, 73]. Thus even though they are not alike, they can still be used as proxies of the “flow”.
@@
We have discussed the denominator of the two efficiencies: $\eta$ and $\theta$. We now discuss the numerator. For what concerns the Portegys algorithm, we have observed that there are more descriptive network structures per se. The accuracy is not necessarily a monotonic decreasing function of the control parameter, $r$. It can exhibit higher values at specific $r$-values as illustrated by the cases of the CIFAR10 and the FashionMNIST.